4. A person stands in an elevator. The person has a mass of 6.0 x 101 kg. Calculate the apparent weight of the person when the elevator is accelerating downwards at 2.5 m/s2


1
Expert's answer
2021-07-20T15:27:43-0400

The apparent weight of the person

"P=m(g-a)=60\:\rm kg\times(9.8-2.5)=438\: N"